The roots of a quadratic are -7 and -2. What is the equation of the quadratic function?
Found 2 solutions by solver91311, stanbon:
Answer by solver91311(24713)



If is a root of a polynomial equation, then is a factor of the polynomial.

John

Egw to Beta kai to Sigma
My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by stanbon(75887)

The roots of a quadratic are -7 and -2. What is the equation of the quadratic function?
---------
If a is a root, x-a is a factor.
-------
f(x) = (x--7)(x--2) = (x+7)(x+2) = x^2 + 9x + 14
===============================================================
Cheers,
Stan H.
============